Robert is four years older than Janet. If the sum of their age is 46, how old are Robert and Janet?
Answer by Boreal(15235)

robert is x
janet is x-4
x+x-4=46
2x-4=46
2x=50
x=25 years, Robert's age
x-4=21 years, Janet's age